If the demand function for a product is given by p=4400/q+3 ; find the elasticity for this demand  function when p = $220. Round your answer off to 2 decimal places. Elasticity = E =
Answer #1

Elasticity = Percentage change in quantity demanded / Percentage change in price

Q1, when price = 1

1 = 4400 / q + 3

q + 3 = 4400 / 1

q = 4400 -3

q = 4397

Q2, when price = 220

220 = 4400 / q+3

q+3 = 4400 / 220

q = 20 - 3

q = 17

Q2 = 17

P1 = 1

P2 = $220

Q1 = 4397

Q2 = 17

Elasticity = Q2 - Q1 / Q1 * P1 / P2 - P1

= 17 - 4397 / 4397 * 1 / 220 - 1

= -4380 / 4397 * 1 / 219

= .997 * .005

= .005

Elasticity = .005